OpenStudy (anonymous):

Which equation has 1 solution?

OpenStudy (solomonzelman):

Typically equations that have one degree. (the biggest power is 1)

OpenStudy (shamil98):

An example would be. x + 3 = 0 The ONLY value of x that makes this equation true is when x = -3. -3 + 3 = 0